OpenStudy (anonymous):

In a college biology class, only half of the class earned a passing grade. Out of all the students, 45% studied and passed. What is the probability of someone having studied, given that they passed?

OpenStudy (anonymous):

is it .45/.5 = 90%

OpenStudy (anonymous):

P(having studies given that they passed) = P(studied and pass) / P(passed) = 0.45/ 0.5 = 0.9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np.